OpenStudy (anonymous):

find the general solution equation for: y'+2y=2t(e^2t) with initial condition: y(0)=1

OpenStudy (amistre64):

you have a characteristic equation of the homogenous part as: r+2=0; therefore r=-2\[y_h=c~e^{-2t}\]

OpenStudy (anonymous):

well my professor said to use integration factor

OpenStudy (amistre64):

or you can use e^(2t) as an integrating thingy

OpenStudy (anonymous):

haha ya i used that and got \[e ^{2y}t=\int\limits e ^{2t} (2te ^{2t})dt\]

OpenStudy (anonymous):

left side is supposed to be with y not t **

OpenStudy (amistre64):

\[\int e^{2t}(y'+2y=2t~e^{2t})~dt \] \[ e^{2t}y=2\int t~e^{4t}~dt \]

OpenStudy (anonymous):

ya i got that but im not sure if i have to use integration by parts or not

OpenStudy (amistre64):

yes, you need to. unless you have remember this setup from the tables ....

OpenStudy (amistre64):

\[\int x~e^{ax}dx=e^{ax}(\frac{1}{a}x-\frac{1}{a^2})\] should be right

OpenStudy (anonymous):

ok so it would be \[=\frac{t ^{4} }{ 2 }e ^{2t}-e ^{2t}/4\]

OpenStudy (amistre64):

umm, no

OpenStudy (amistre64):

close tho

OpenStudy (amistre64):

\[e^{2t}y=e^{4t}(\frac t{4}-\frac{1}{16})+C\]

OpenStudy (anonymous):

o lol i see where i messed up

OpenStudy (amistre64):

:) then of course its left with algebra, divide off the e^{2t} and insert your initial condition to determine a suitable C

OpenStudy (anonymous):

yep got it

OpenStudy (amistre64):

good job ... you find out that its not so much the calculus that gets you; its all the little mistakes we tend to do in the algebra that kills us lol